A cellular phone company offers a contract for which the cost C, in dollars, of t minutes of telephoning is given by C=0.25(t-500)+47.95, where it is assumed that t=>500 minutes. What times will keep costs between $89.45 and $132.70?
(Round to the nearest whole number)

Answer by nerdybill(7384)

A cellular phone company offers a contract for which the cost C, in dollars, of t minutes of telephoning is given by C=0.25(t-500)+47.95, where it is assumed that t=>500 minutes. What times will keep costs between $89.45 and $132.70?
(Round to the nearest whole number)
.
Minutes when $89.45:
89.45=0.25(t-500)+47.95
41.50=0.25(t-500)
41.50=0.25t-125
166.5=0.25t
166.5/0.25=t
666 minutes = t
.
Minutes when $132.70:
132.70=0.25(t-500)+47.95
84.75=0.25(t-500)
84.75=0.25t-125
209.75=0.25t
209.75/0.25=t
839 minutes = t
.
answer:
666 <= t <= 839